# A Universal Scheme for Dynamic Partitioned Shortest Path Index

### v2

@@ 5. Universal PSP Index Scheme and Newly Generated PSP Indexes
In this section, we propose a universal PSP index scheme and generator and put forward five representative PSP indexes for query-oriented or update-oriented applications based on this scheme.
@@
5.1. PSP Index Scheme and Generator
@@
After elaborating on each component of the PSP index, we now have our PSP index scheme ready as shown in Figure 9. The main part is the PSP index generator, which is formed by coupling those three components. For easy selection of the PSP strategy and SP algorithm, we sort all of their choices roughly in terms of query and update efficiency. As shown in Figure 9, the lighter color signifies faster efficiency for PSP strategy and SP algorithm. Meanwhile, planar and hierarchical partitions are more suitable for low-treewidth networks, while core-periphery can deal with high-treewidth graphs. Given the specific application requirements in terms of computation efficiency, space consumption, or partition type, our scheme could provide one PSP index with a specified PSP strategy, SP algorithm, and partition structure through the PSP index generator. For starters, the existing PSP indexes can find their positions in our scheme as will be discussed in Section 7, and we can compare them theoretically and fairly. Apart from this, we can further utilize it as a PSP Index Generator to create new PSP indexes that are suitable to user requirements, especially when all the existing ones are query-oriented, while there leaves a vast space of combinations for new ones. Specifically, the generated PSP index needs to choose the components from each of the three dimensions, and its name also has three parts corresponding to the three dimensions. Although we can enumerate them all, we choose to introduce the representative ones from the perspectives of query- and update-oriented under different partition structures, as shown in Figure 9.
@@ 5.2. Representative Generated New PSP Indexes
1. Query-Oriented Planar PSP Index: P-TD-P. Given a planar partition, the PSP strategies can apply directly to it. Next, we introduce how to construct the PSP indexes that are efficient in query processing under planar partition. Firstly in terms of PSP strategy, since both Pre-Boundary and Post-Boundary are efficient in query answering and our proposed Post-Boundary is faster than Pre-Boundary in index construction, we use it as the PSP strategy. Secondly, in terms of the SP algorithm, we could choose TD as the index for both partitions and overlay graph. Although all-pair is faster, its space consumption is intolerable. Its structure is shown in Figure 9-(a). Specifically, the index construction takes $O(V_{max}\cdot(\log V_{max}+h_{max}\cdot w_{max}))$ time for the partition TD, $O(V_{\tilde{G}}(\log V_{\tilde{G}}+h_{\tilde{G}}\cdot w_{\tilde{G}}))$ for the overlay TD, $O(B^{2}_{max}w_{\tilde{G}})$ for boundary correction, and $O(w_{\tilde{G}}+w^{2}_{max}\cdot\delta)$ for the partition TD refresh, where $\delta$ is the affected shortcut number and $max$ are the corresponding max values in the partitions. For the query processing, the intra-query takes $O(w_{max})$ as partition index is correct, and the inter-query takes $O(max\{B_{max}w_{max},B^{2}{max}w_{\tilde{G}}\})$ for the partition and overlay query and $O(B^{2}_{max})$ for the combinations. For the index update, it takes $O(w^{2}_{max}\delta)$ to update partition TD, $O(w^{2}_{\tilde{G}}\delta)$ to update overlay TD, and $O(B^{2}_{max}w_{\tilde{G}})$ to check boundaries.
@@
3. Query-Oriented Core-Tree PSP Index: P-PT-CP. The core-periphery partition index comprises the core index $L_{c}$ and the periphery index $\{L_{i}\}$. It seems that $L_{c}$ and $L_{i}$ can be constructed by their corresponding subgraphs $G_{c}$ and $G_{i}$, respectively. Although the core does not belong to any partition, they are connected to the partitions, and we treat the core as the overlay graph. Different from the previous planar partition, the core here usually has a large degree so its index is limited to PLL. As for sketch, we omit it here because its PLL core + pruned direct search seems to be the only solution for huge networks. Next, we discuss the remaining parts. For the PSP strategy, Post-Boundary is utilized as the queries within the periphery can be handled without the core index. The periphery index uses TD because the periphery usually has a small degree. This structure is shown in Figure 9-(c). In terms of index construction, because the periphery is constructed through contraction, we regard it as a by-product of the partition phase and do not construct their labels in the first step. Then it takes $O(w_{c}E_{c}\log V_{c}+w^{2}_{c}V_{c}\log^{3}V_{c})$ for the core PLL, $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction, and $O(V_{max}\cdot w^{2}_{max}\cdot\log V_{max})$ for periphery label. In terms of query processing, the intra-query takes $O(w_{max})$ time. The inter-query takes $O(max\{B_{max}w_{max},B^{2}_{max}w_{c}\log V_{c}\})$ for the periphery and core, and $O(B^{2}_{max})$ for the combinations. In terms of index update, it takes $O(w^{2}_{max}\cdot\delta)$ for the periphery TD, $O(w_{c}E_{c}\log V_{c})$ for the core PLL, and $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction.
@@
Remark 3. We note that these five PSP indexes are novel structures, each with unique technical challenges, especially in updating. Nevertheless, we believe the PSP scheme holds greater significance than these new indexes, as it can guide new PSP index designs. We omit their details due to limited space, but their source codes are provided (GitHub, [n.d.]). Besides, we take the N-CH-P index as an example and elaborate on its details in the Appendix (FullVersion).
@@
[] Type Name Dataset $|V|$ $|E|$ Road Network NY 1 New York City 264,346 730,100 FL 1 Florida 1,070,376 2,687,902 W 1 Western USA 6,262,104 15,119,284 US 1 Full USA 23,947,347 57,708,624 Complex Network GO 2 Google 855,802 8,582,704 SK 3 Skitter 1,689,805 21,987,076 WI 3 Wiki-pedia 3,333,272 200,923,676 FR 2 com-Friendster 65,608,366 3,612,134,270 [1] http://www.dis.uniroma1.it/challenge9/download.shtml [2] http://snap.stanford.edu/data [3] http://konect.cc/
Table 3. Average Vertex Degree of Overlay Graph
Dataset NY FL GO SK
Method PUNCH HEP PUNCH HEP KaHyPar HEP KaHyPar HEP
w/o Opt 48.68 325.98 36.17 719.25 947.36 9754.52 1420.18 34889.5
w/ Opt 48.68 320.61 36.17 714.30 912.24 7678.05 1409.98 27326.5
6. Experimental Evaluation
In this section, we evaluate the effectiveness of the proposed methods and provide a systematic evaluation of PSP indexes. All the algorithms are implemented in C++ with full optimization on a server with 4 Xeon Gold 6248 2.6GHz CPUs (total 80 cores / 160 threads) and 1.5TB memory. The thread number is set to 150.
@@ 6.1. Experimental Settings
Datasets and Queries. We test on eight real-life datasets, including four weighted road networks (NY, FL, W, and US) and four complex networks (Web graph: GO and WI; Social network: SK and FR), as shown in Table 2. The edge weight of complex networks is generated by following the manner (Zhang et al., 2021b) that is inversely proportional to the highest degree of the endpoints. We randomly generate 10,000 queries and 1,000 update instances for each dataset to assess the query processing and index maintenance efficiency, respectively.
@@
Partition Methods and Their Performance Metrics. To select the suitable partition method for PSP index, we implement 8 representative partition methods: Bubble (Diekmann et al., 2000), RCB (Simon, 1991), KaHyPar (Akhremtsev et al., 2017), PUNCH (Delling et al., 2011), SCOTCH (Barat et al., 2018), METIS (Karypis and Kumar, 1998), HEP (Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022), all of which are planar partition methods with PUNCH and RCB only working on road networks. Since core-periphery partition (including Tree-Decomposition-based partition (Ouyang et al., 2018) and landmark selection in Sketch (Farhan et al., 2022)) and hierarchical partition (mostly using METIS (Karypis and Kumar, 1998)) methods are finite, they are selected by default in the corresponding PSP index. We evaluate the performance of a partition method from three aspects: overall border number $|B|=\sum_{i\in[1,k]}|B_{i}|$, average partition border number $\overline{|B_{i}|}=\frac{\sum_{i\in[1,k]}|B_{i}|}{k}$, and partition connectivity $R_{C}=\frac{\sum_{i\in[1,k]}|C(G_{i})|}{k}$, where $C(G_{i})$ is the connected component number of subgraph $G_{i}$.
@@
PSP Indexes and Their Performance Metrics In terms of PSP indexes comparison, we evaluate 5 existing representative PSP indexes spreading over our proposed three partition structures: a) FHL (Liu et al., 2021, 2022a) and THop (Li et al., 2019c) belongs to planar partition; b) Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) and Sketch (Farhan et al., 2022) belong to the core-periphery partition; c) G-Tree (Zhong et al., 2013) belongs to the hierarchical partition. Besides, we implement our 5 newly proposed PSP indexes (see Section 5): P-TD-P, N-CH-P, P-PT-CP, N-PC-CP, and N-TS-HP. For the state-of-the-art shortest path algorithms CH (Geisberger et al., 2008; Ouyang et al., 2020), TD (Ouyang et al., 2018; Zhang et al., 2021c), and PLL (Akiba et al., 2013; Zhang et al., 2021b), we implement their partitioned version (named PCH, PTD, and PPLL respectively) to test the performance of different PSP strategies and select the suitable partition number. We measure the performance of the PSP index from four aspects: index construction time $t_{c}$, query time $t_{q}$, update efficiency $t_{u}$, and index size $s$. Note that we do not report the result (or set it to $\infty$) if the running time of a partition method or PSP index exceeds 24 hours.
@@
Parameter Setting. We set the default partition number as $32$ for all planar partitions and the bandwidth of core-tree indexes as $40$ as per experimental results. The landmark number of Sketch is set to $20$ by referring to (Farhan et al., 2022). As for the hierarchical partition, we follow G-tree (Zhong et al., 2013; Li et al., 2019a) and set the fan-out $f$ as $4$ and the maximum leaf node size $\tau$ as 128 in NY, 256 in FL, 512 in W and US, respectively.
Figure 10. Partition Performance When Varying $k$.
@@
Figure 12. Effect of Partition Number. Bar: Left, Ball: Right
Figure 13. Comparison with Non-partitioned SP Indexes. Bar: Left, Ball: Right
Figure 14. Effectiveness of Partitioned Shortest Path Strategy. (Bar: Left Axis; Ball: Right Axis)
@@ 6.2. Experiment Result Analysis
Exp 1: Performance of Partition Methods When Varying $k$. We first report the performance of different partition methods with varying partition number $k$ (we only report results on FL, US, and GO due to limited space; the full results are in the full version (Zhang et al., 2023b))). As shown in Figure 10 (a)-(f), the overall border number $|B|$ generally increases, while the average border number $\overline{|B_{i}|}$ decreases as $k$ grows from 4 to 256. This indicates that having a larger partition number can potentially speed up the query processing of the PSP index due to a smaller average partition border number. However, it also leads to a larger overlay graph, which can slow down the query processing and construction of the overlay index. Therefore, there should be a trade-off between them, and the best partition number typically varies for different PSP indexes. Furthermore, as per Figure 10 (g)-(i), only PUNCH and Bubble have a partition connectivity ratio $R_{C}$ of 1, which implies that at least one partition generated by other partition methods is unconnected (with multiple separate components). Since PTD requires the vertices within one partition to be connected, only PUNCH and Bubble suit it.

### v3

@@ 6. Universal PSP Index Scheme and Newly Generated PSP Indexes
In this section, we propose a universal PSP index scheme and then put forward five representative PSP indexes for query-oriented or update-oriented applications based on this scheme.
@@
6.1. PSP Index Dimensions and Universal Scheme
Based on the comprehensive literature review and analysis of the PSP index in previous Sections, we identify three critical dimensions of the PSP index as follows.
1) Shortest Path Algorithm that decides the underlying methods for the overlay graph and partitions. The performance of different SP algorithms on dynamic networks has been thoroughly studied in our previous work (Zhang et al., 2021a);
2) Graph Partition Method, which divides the graph into multiple subgraphs and is the key difference from the standalone SP indexes. The existing partition methods are not designed for path indexes, so how they affect the PSP index is unknown and deserves study;
3) PSP Strategy, which provides a procedure to organize the index construction and update for different subgraphs and coordinate those indexes such that the query correctness is guaranteed. This dimension is crucial because it determines the complexity of both the query and update. However, only straightforward pre-boundary strategy exists currently. Thus we further propose two novel strategies to improve its index construction and maintenance performance.
@@ 6.1. PSP Index Dimensions and Universal Scheme
Given a dynamic network, these three dimensions provide a roadmap to construct a PSP index by selecting a specific graph partition method, SP algorithm and PSP strategy such that the application requirement (query / update efficiency) can be satisfied. Inspired by this, we propose a universal PSP Index Scheme as shown in Figure 9. The main part is the PSP index generator, which is formed by coupling those three components. For easy selection of the PSP strategy and SP algorithm, we sort all of their choices roughly in terms of query and update efficiency. As shown in Figure 9, the lighter color signifies faster efficiency for PSP strategy and SP algorithm. Meanwhile, planar and hierarchical partitions are more suitable for low-treewidth networks, while core-periphery can deal with high-treewidth graphs.
There are two benefits of our PSP index scheme. Firstly, the existing PSP indexes can find their positions in our scheme and we can compare them theoretically and fairly. For exmaple, 1) Pre-Boundary + Search/All-Pair + Hierarchy/ Planar: HiTi (Jung and Pramanik, 2002), CRP (Delling et al., 2017), PbS (Chondrogiannis and Gamper, 2014), ParDiSP (Chondrogiannis and Gamper, 2016) pre-compute shortest distances between boundaries to guide the search. In a broad sense, Arc-flag (Möhring et al., 2007) and SHARC (Bauer and Delling, 2010) also belong to this category. G-tree (Zhong et al., 2013; Li et al., 2019a) uses dynamic programming to compute the distance between layer’s all-pair information to replace searching and it is used in kNN (Zhong et al., 2015), ride-sharing (Ta et al., 2017), time-dependent (Wang et al., 2019), and distance embedding (Huang et al., 2021). They are slow to construct due to Pre-B and inefficient to query due to direct search; 2) Pre-Boundary+Search/PLL+Planar: COLA (Wang et al., 2016) builds labels for the skyline shortest path on the overlay graph for the constrained shortest path. Its construction suffers from Pre-B and query suffers from searching; 3) Pre-Boundary+TD/TD+Planar: FHL (Liu et al., 2021, 2022) builds the TD both within and between partitions for multi-dimensional skyline paths; 4) Pre-Boundary+PLL/PLL+Planar: T2Hop (Li et al., 2019c, 2020b) utilizes two layers of PLL to reduce the complexity of long-range time-dependent paths, and this structure’s performance is limited by PLL; 5) Pre-Boundary+Search/All-Pair/PLL+Sketch: This category works on huge graphs where index is impossible so only a small number of landmarks are selected to prune the search (Das Sarma et al., 2010; Wang et al., 2021; Farhan et al., 2022) or approximate the result (Gubichev et al., 2010; Qiao et al., 2012).
Secondly, given the specific application requirements in terms of computation efficiency, space consumption, or partition type, our scheme could provide one PSP index with a specified PSP strategy, SP algorithm, and partition structure through the PSP index generator. Especially when all the existing ones are query-oriented, there is a vast space of combinations for new ones. Specifically, the generated PSP index needs to choose the components from each of the three dimensions, and its name also has three parts corresponding to the three dimensions. Although we can enumerate them all, we choose to introduce five representative ones from the perspectives of query- and update-oriented under different partition structures, as shown in Figure 9. The details of these five PSP indexes will be elaborated in Section 6.2.
@@ 6.2. Representative Generated New PSP Indexes
1. Query-Oriented Planar PSP Index: P-TD-P. We introduce how to construct a planar PSP index that is efficient in query processing. Firstly, in terms of PSP strategy, since both Pre-Boundary and Post-Boundary are efficient in query answering and our proposed Post-Boundary is generally faster than Pre-Boundary in index maintenance, we use it as the PSP strategy. Secondly, in terms of the SP algorithm, we could choose TD as the index for both partitions and overlay graph. Although all-pair is faster, its space consumption is intolerable. Its structure is shown in Figure 9-(a). Specifically, the index construction takes $O(V_{max}\cdot(\log V_{max}+h_{max}\cdot w_{max}))$ time for the partition TD, $O(V_{\tilde{G}}(\log V_{\tilde{G}}+h_{\tilde{G}}\cdot w_{\tilde{G}}))$ for the overlay TD, $O(B^{2}_{max}w_{\tilde{G}})$ for boundary correction, and $O(w_{\tilde{G}}+w^{2}_{max}\cdot\delta)$ for the partition TD refresh, where $\delta$ is the affected shortcut number and $max$ are the corresponding max values in the partitions. For the query processing, the intra-query takes $O(w_{max})$ as partition index is correct, and the inter-query takes $O(max\{B_{max}w_{max},B^{2}{max}w_{\tilde{G}}\})$ for the partition and overlay query and $O(B^{2}_{max})$ for the combinations. For the index update, it takes $O(w^{2}_{max}\delta)$ to update partition TD, $O(w^{2}_{\tilde{G}}\delta)$ to update overlay TD, and $O(B^{2}_{max}w_{\tilde{G}})$ to check boundaries.
@@
3. Query-Oriented Core-Tree PSP Index: P-PT-CP. The core-periphery partition index comprises the core index $L_{c}$ and the periphery index $\{L_{i}\}$. Although the core does not belong to any partition, they are connected to the partitions, and we treat the core as the overlay graph. Different from the previous planar partition, the core here usually has a large degree so its index is limited to PLL. As for sketch, we omit it here because its PLL core + pruned direct search seems to be the only solution for huge networks. Next, we discuss the remaining parts. For the PSP strategy, Post-Boundary is utilized as the queries within the periphery can be handled without the core index. The periphery index uses TD because the periphery usually has a small degree. This structure is shown in Figure 9-(c). In terms of index construction, because the periphery is constructed through contraction, we regard it as a by-product of the partition phase and do not construct their labels in the first step. Then it takes $O(w_{c}E_{c}\log V_{c}+w^{2}_{c}V_{c}\log^{3}V_{c})$ for the core PLL, $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction, and $O(V_{max}\cdot w^{2}_{max}\cdot\log V_{max})$ for periphery label. In terms of query processing, the intra-query takes $O(w_{max})$ time. The inter-query takes $O(max\{B_{max}w_{max},B^{2}_{max}w_{c}\log V_{c}\})$ for the periphery and core, and $O(B^{2}_{max})$ for the combinations. In terms of index update, it takes $O(w^{2}_{max}\cdot\delta)$ for the periphery TD, $O(w_{c}E_{c}\log V_{c})$ for the core PLL, and $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction.
@@
In summary, these five PSP indexes are novel structures, each with unique technical challenges, especially in updating. Nevertheless, we believe the PSP scheme holds greater significance than these new indexes, as it can guide new PSP index designs. We omit their details due to limited space, but their source codes are provided (GitHub, [n. d.]). Besides, we take the N-CH-P index as an example and elaborate on its details in the extended version (Zhang et al., [n. d.]).
7. Experimental Evaluation
In this section, we evaluate the effectiveness of the proposed methods and provide a systematic evaluation of PSP indexes. All the algorithms are implemented in C++ with full optimization on a server with 4 Xeon Gold 6248 2.6GHz CPUs (total 80 cores / 160 threads) and 1.5TB memory. The default thread number is set to 150.
@@ 7. Experimental Evaluation
[] Category Name Dataset $|V|$ $|E|$ Type Road Network NY 1 New York City 264,346 730,100 road FL 1 Florida 1,070,376 2,687,902 road W 1 Western USA 6,262,104 15,119,284 road US 1 Full USA 23,947,347 57,708,624 road Complex Network GO 2 Google 855,802 8,582,704 web SK 3 Skitter 1,689,805 21,987,076 social WI 3 Wiki-pedia 3,333,272 200,923,676 web FR 2 com-Friendster 65,608,366 3,612,134,270 social [1] http://www.dis.uniroma1.it/challenge9/download.shtml
[2] http://snap.stanford.edu/data [3] http://konect.cc/
@@ 7.1. Experimental Settings
Datasets. We test on eight real-life datasets (four road networks and four complex networks as shown in Table 3). Following (Zhang et al., 2021b; Zhou et al., 2024b; Zhang et al., 2021a; Zhou et al., 2024a), the edge weights of complex networks are randomly generated inversely proportional to the highest degree of the endpoints.
@@
Partition Methods and Their Performance Metrics. To select the suitable partition method for PSP index, we implement 8 representative methods: Bubble (Diekmann et al., 2000), RCB (Simon, 1991), KaHyPar (Akhremtsev et al., 2017), PUNCH (Delling et al., 2011b), SCOTCH (Barat et al., 2018), METIS (Karypis and Kumar, 1998), HEP (Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022), all of which are planar with PUNCH and RCB only working on road networks. Since core-periphery (core-tree (Zhou et al., 2024b) and Sketch (Farhan et al., 2022)) and hierarchical partition (uses METIS (Karypis and Kumar, 1998)) methods are finite, they are selected by default in the corresponding PSP index. We evaluate the performance of a partition method from three aspects: overall border number $|B|=\sum_{i\in[1,k]}|B_{i}|$, average partition border number $\overline{|B_{i}|}=\frac{\sum_{i\in[1,k]}|B_{i}|}{k}$, and partition connectivity $R_{C}=\frac{\sum_{i\in[1,k]}|C(G_{i})|}{k}$, where $C(G_{i})$ is the connected component number of subgraph $G_{i}$.
@@
PSP Indexes and Performance Metrics. We evaluate 5 SOTA PSP indexes combining with the three proposed partition structures: a) FHL (Liu et al., 2021, 2022) and THop (Li et al., 2019c) from planar partition; b) Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022; Zhou et al., 2024b) and Sketch (Farhan et al., 2022) from core-periphery; c) G-Tree (Zhong et al., 2013) from hierarchical partition. Besides, we implement our 5 newly proposed PSP indexes: P-TD-P, N-CH-P, P-PT-CP, N-PC-CP, and N-TS-HP. For the SOTA shortest path indexes CH (Geisberger et al., 2008; Ouyang et al., 2020), TD (Ouyang et al., 2018; Zhang et al., 2021c), and PLL (Akiba et al., 2013; Zhang et al., 2021b), we implement their partitioned version (named PCH, PTD, and PPLL). We measure the performance from four aspects: index construction time $t_{c}$, query time $t_{q}$, update time $t_{u}$, and index size $s$. We do not report the result ($\infty$) if index time exceeds 24 hours.
Query and Update. For each dataset, we randomly generate a set of 10,000 queries and 1,000 edge update instances to evaluate $t_{q}$ and $t_{u}$. Following (Zhang et al., 2021c, a), for each selected edge $e$, we decrease or increase its weight to $\alpha\cdot|e|$, where $\alpha$ randomly ranges in $(0,2]$. Besides, we also randomly generate 1000 update instances for each type of structural update (edge/vertex insertion/deletion) for Exp 5.
@@
Parameter Setting. We set the default partition number as $32$ for all planar PSP and the bandwidth of CT-based PSP as $40$ as per our preliminary experimental results in (Zhang et al., [n. d.]). The landmark number of Sketch is $20$ (Farhan et al., 2022). For the hierarchical PSP, we follow G-tree (Zhong et al., 2013; Li et al., 2019a) and set the fan-out $f$ as $4$ and the leaf node size $\tau$ as 128 in NY, 256 in FL, 512 in W and US. We use No-boundary as the default strategy if the PSP strategy is not specified.
Figure 10. Performance of Partition Methods When Varying Partition Number $k$.
@@ 7.2. Experiment Result Analysis
Exp 1: Performance of Partition Methods When Varying $k$. As shown in Figure 10 (a)-(p), the overall border number $|B|$ generally increases while the average border number $\overline{|B_{i}|}$ decreases with the increase of $k$. This indicates that having a larger $k$ can potentially speed up the query processing of the partition index due to a smaller average partition border number. However, it also leads to a larger overlay graph, which can slow down the query processing and construction of the overlay index. Therefore, there is a trade-off between them. Furthermore, as per Figure 10 (q)-(x), among the partition methods used, only PUNCH and Bubble have a partition connectivity ratio $R_{C}$ of 1. This implies that the partitions generated by other partition methods generally contain multiple unconnected components. Since PTD requires the vertices within the same partition to be connected, only PUNCH and Bubble suit it. Overall, PUNCH and KaHyPar perform best in $|B|$ and $\overline{|B_{i}|}$ on road networks. When dealing with complex networks, it is advisable to use KaHyPar, METIS, or HEP. In the case of the largest graph FR, only HEP and CLUGP are capable of completing graph partitioning within 24 hours.